### v1

@@ Abstract
We present the analysis of over 3,000 red-Herschel sources () using public data from the ALMA archive and the Herschel-ATLAS survey. This represents the largest sample of red-Herschel sources with interferometric follow-up observations to date. The high ALMA angular resolution and sensitivity (1 arcsecond; mJy beam-1) allow us to classify the sample into individual sources, multiple systems, and potential lenses and/or close mergers. Interestingly, even at this high angular resolution, 73 per cent of our detections are single systems, suggesting that most of these galaxies are isolated and/or post-merger galaxies. For the remaining detections, 20 per cent are classified as multiple systems, 5 per cent as lenses and/or mergers, and 2 per cent as low- galaxies or Active Galactic Nuclei. Combining the Herschel/SPIRE and ALMA photometry, these galaxies are found to be extreme and massive systems with a median star formation rate of 1,500 and molecular gas mass of . The median redshift of individual sources is , while the likely lensed systems are at , with redshift distributions extending to . The inferred depletion times are consistent with the starburst-type population with a mild redshift evolution. Our results suggest a common star-formation mode for extreme galaxies across cosmic time, likely triggered by close interactions or disk-instabilities. Moreover, all galaxies with mJy are gravitationally amplified which, similar to the established mJy threshold, can be used as a simple criterion to identify gravitationally lensed galaxies.
@@
The SMG are characterized by having infrared luminosities () of L⊙, equivalent to those of Ultra-luminous infrared galaxies (ULIRGs) and even reaching the extreme luminosities of Hyper-luminous infrared galaxies (HyLIRGs, L⊙). These luminosities correspond to star formation rates (SFR) of M⊙ yr-1. Some spectroscopic studies have probed the CO line emission in this population, finding, typically, large reservoirs of molecular gas () distributed in a varied morphology that includes disk galaxies and compact star formation bursts of gas-rich galaxies in interaction (e.g. Oteo et al., 2016). The inferred depletion times from these observations are at the order of tens of Myrs, which is short compared to the 1 Gyr depletion time in normal galaxies (e.g. Tacconi et al., 2010; Saintonge et al., 2011). This difference opens the question of what is triggering the SFR in these galaxies, the high content of molecular gas or physical processes that are different in the early universe (see reviews by Blain et al., 2002; Carilli & Walter, 2013; Casey et al., 2014).
@@
This paper is structured as follows: Section 2 provides a detailed description of the sample selection, data processing and source detection; Section 3 presents the methodology and results; Section 4 compares our results in the framework of the different studies of red-Herschel galaxies as in the general scheme of the SMG population; finally, we summarize our results and conclusions in Section 5.
@@ 2.1 Red H-ATLAS galaxies observed with ALMA
We build a sample of red Herschel galaxies (defined with a simple selection criterion of ) using the data release catalogs from H-ATLAS (Valiante et al., 2016; Maddox et al., 2018). We find 6,194 sources that meet the criterion, representing only 1.4 per cent of all the H-ATLAS detections.
@@ 3.1 Classification
Fields with at least one source detected above the adopted threshold (i.e. 2416 fields) were classified in three different groups depending on the number of detections and their proximity. Examples of the three classifications, which are described in detail below, can be found in Fig. 4.
@@ 3.1.3 M - Fields
Multiple fields include those with two or more sources that are separated by more than 3 arcseconds. This category contains 475 systems. From these, 421 have only two components, 51 have three sources, and 3 fields have 4 detections. The sources in these fields have an average flux density of 2.7 mJy at 1.3 mm. The analysis and physical properties of these fields will be presented in a forthcoming paper.
@@ 4.2 Multiplicity
Working with samples of similar galaxies (although significantly more limited in size), studies such as Ma et al. (2019), Greenslade et al. (2020), Montaña et al. (2021), and Cox et al. (2023) reported that the multiplicity of the red-Herschel sources ranges between 9 and 27 per cent, which is in good agreement with our results. Greenslade et al. (2020) and Montaña et al. (2021) further suggested that, if their non-detections were multiple systems with individual components below the detection limit of the higher angular resolution observations, the multiplicity fraction might increase to per cent. If we assume that the red-Herschel sources in our sample undetected in the ALMA 1.3 mm maps are multiple systems, the multiplicity fraction would reach per cent.
@@ 4.6 Depletion times and Star formation efficiencies
The depletion times (), which are independent of gravitational amplification to a first order, can give clues about the processes behind the extreme SFR in these galaxies. We use the values estimated in Sections 3.2 and 3.3 to determine depletion times (and star formation efficiencies (SFE), i.e. ) for the sources in our sample. These results, however, should be taken with caution since the are inferred from the ALMA 1.3 mm photometry and are not totally independent from the SED-based SFR determinations.
The inferred depletion times for the S - Fields cover a range from 0.2 Gyr to 0.4 Gyr. These values are consistent with those found in the literature for similar samples, where the depletion times are between 0.1 Gyr and 1 Gyr (Hatsukade et al., 2015; Pantoni et al., 2021; Berta et al., 2023; Castillo et al., 2023).
On the other hand, our estimated depletion times are lower compared to those found for nearby galaxies ( and Gyr, e.g Bigiel et al., 2008; Leroy et al., 2008; Genzel et al., 2010). If we scale our gas mass determinations to assume , our depletion times are reduced even further (30 to 60 Myr). These values are in agreement with the extreme SMG population reported to have high star formation efficiency (Aravena et al., 2016; Yang et al., 2017; Ikarashi et al., 2022).
Figure 12 shows the depletion times estimated for our samples versus redshift, along with the expected evolution for main-sequence star-forming galaxies and starburst type galaxies (Tacconi et al., 2018). The depletion times of our samples are systematically shorter than those expected for main-sequence galaxies and consistent with the starburst population, showing a mild redshift evolution towards shorter times at higher redshifts.
@@ 5 Summary and Conclusions
The median IR luminosities for the S - Fields and PLUM - Fields of and L⊙, correspond to SFR of and , respectively. This is in line with previous studies and confirms the extreme nature of these SMG, which also show high gas and dust content with and . Combining these measurements, we found average depletion times of Gyr, consistent with what has been measured for other SMG, implying high SFE compared to the main-sequence population. Interestingly, the inferred depletion times (and, thus, the SFE) show a mild redshift evolution, consistent with previous result in the literature. We stress, however, that these results should be taken with caution until a more direct measurement of the gas masses become available.
@@ Acknowledgements
We thank to the anonymous referee for insightful comments that have improved the manuscript. We also want to thank Olga Vega, Yalia Divakara Mayya, Manuel Zamora-Avilés, Elena Terlevich and Eric F. Jiménez-Andrade for their helpful discussions and comments, and to Michał Michałowski for insightful suggestions on the flux-deblending process. MQR would like to thank the Consejo Nacional de Humanidades Ciencias y Technologías (CONAHCYT) for her PhD grant. This work has been supported by the CONAHCYT through projects A1-S-45680 and CB 2016 - 281948. JAZ acknowledge funding from JSPS KAKENHI grant number KG23K13150. The Herschel-ATLAS is a project from Herschel, which is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. This paper use the following ALMA projects: ADS/JAO.ALMA#2016.1.00087.S, ADS/JAO.ALMA#2016.1.00139.S, ADS/JAO.ALMA#2017.1.00510.S, ADS/JAO.ALMA#2018.1.00489.S, ADS/JAO.ALMA#2018.1.00526.S, ADS/JAO.ALMA#2021.1.01628.S and ADSJAO.ALMA#2022.1.00432.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.

### v2

@@ Abstract
We present the analysis of over 3,000 red-Herschel sources () using public data from the ALMA archive and the Herschel-ATLAS survey. This represents the largest sample of red-Herschel sources with interferometric follow-up observations to date. The high ALMA angular resolution and sensitivity (1 arcsecond; mJy beam-1) allow us to classify the sample into individual sources, multiple systems, and potential lenses and/or close mergers. Interestingly, even at this high angular resolution, 73 per cent of our detections are single systems, suggesting that most of these galaxies are isolated and/or post-merger galaxies. For the remaining detections, 20 per cent are classified as multiple systems, 5 per cent as lenses and/or mergers, and 2 per cent as low- galaxies or Active Galactic Nuclei. Combining the Herschel/SPIRE and ALMA photometry, these galaxies are found to be extreme and massive systems with a median star formation rate of 1,500 and molecular gas mass of . The median redshift of individual sources is , while the likely lensed systems are at , with redshift distributions extending to . Our results suggest a common star-formation mode for extreme galaxies across cosmic time, likely triggered by close interactions or disk-instabilities, and with short depletion times consistent with the starburst-type population. Moreover, all galaxies with mJy are gravitationally amplified which, similar to the established mJy threshold, can be used as a simple criterion to identify gravitationally lensed galaxies.
@@
The SMG are characterized by having infrared luminosities () of L⊙, equivalent to those of Ultra-luminous infrared galaxies (ULIRGs) and even reaching the extreme luminosities of Hyper-luminous infrared galaxies (HyLIRGs, L⊙). These luminosities correspond to star formation rates (SFR) of M⊙ yr-1. Some spectroscopic studies have probed the CO line emission in this population, finding, typically, large reservoirs of molecular gas () distributed in a varied morphology that includes disk galaxies and compact star formation bursts of gas-rich galaxies in interaction (e.g. Oteo et al., 2016). The inferred depletion times from these observations are at the order of tens to a few hundred Myrs, which is short compared to the 1 Gyr depletion time in normal galaxies (e.g. Tacconi et al., 2010; Saintonge et al., 2011). This difference opens the question of what is triggering the SFR in these galaxies, the high content of molecular gas or physical processes that are different in the early universe (see reviews by Blain et al., 2002; Carilli & Walter, 2013; Casey et al., 2014).
@@
This paper is structured as follows: Section 2 provides a detailed description of our sample, the Herschel-Alma Red sources Public Archive Study (HARPAS), as well as of the data processing and source detection; Section 3 presents the methodology and results; Section 4 compares our results in the framework of the different studies of red-Herschel galaxies as in the general scheme of the SMG population; finally, we summarize our results and conclusions in Section 5.
@@ 2.1 Red H-ATLAS galaxies observed with ALMA
We build a sample of red-Herschel galaxies (defined with a simple selection criterion of ) using the data release catalogs from H-ATLAS (Valiante et al., 2016; Maddox et al., 2018). We find 6,194 sources that meet the criterion, representing only 1.4 per cent of all the H-ATLAS detections.
@@ 3.1 Classification
Fields with at least one source detected above the adopted threshold (i.e. 2416 fields) were classified in three different groups depending on the number of detections and their proximity. Examples of the three classifications, which are described in detail below, can be found in Fig. 4. Additionally, fields with no detections are classified as N - Fields and are briefly discussed in Section 3.1.4.
@@ 3.1.3 M - Fields
Multiple fields include those with two or more sources that are separated by more than 3 arcseconds. This category contains 474 systems. From these, 420 have only two components, 51 have three sources, and 3 fields have 4 detections. The sources in these fields have an average flux density of 2.7 mJy at 1.3 mm. The analysis and physical properties of these fields will be presented in a forthcoming paper.
3.1.4 N - Fields
This classification corresponds to fields where no detections above a 5 level are found. There are 673 fields in this category with a median r.m.s. of 0.19 mJy beam-1. The distribution of r.m.s. values in these N - Fields is similar to those from the other categories where sources are detected. A fraction of these fields may therefore correspond to multiple systems of sources with flux densities below the 5 detections threshold, as it has been suggested in previous follow-up studies of red-Herschel sources (e.g. Greenslade et al., 2020; Montaña et al., 2021).
@@ 3.2 Redshifts, Luminosities and Star Formation Rates
In an attempt to independently estimate the SFRs and gas masses (which are based on the ALMA photometry; Sec. 3.3), we run MMPz using only the three Herschel/SPIRE data-points. This alternative, however, resulted in loose constraints on the SFR and, in some cases, overestimated redshifts, highlighting the importance of having photometric information probing the Rayleigh-Jeans regime. We therefore base our analysis and conclusions on the more robust and reliable results derived from the fit to the full dataset (i.e. using the Herschel/SPIRE and 1.3 mm ALMA photometry).
@@ 4.2 Multiplicity
Working with samples of similar galaxies (although significantly more limited in size), studies such as Ma et al. (2019), Greenslade et al. (2020), Montaña et al. (2021), Bendo et al. (2022), and Cox et al. (2023) reported that the multiplicity of the red-Herschel sources ranges between 9 and 27 per cent, which is in good agreement with our results. Greenslade et al. (2020) and Montaña et al. (2021) further suggested that, if their non-detections were multiple systems with individual components below the detection limit of the higher angular resolution observations, the multiplicity fraction might increase to per cent. If we assume that the red-Herschel sources in our sample undetected in the ALMA 1.3 mm maps are multiple systems, the multiplicity fraction would reach per cent.
@@ 4.6 Depletion times and Star formation efficiencies
The depletion times (), which are independent of gravitational amplification to a first order, can give clues about the processes behind the extreme SFR in these galaxies. We use the values estimated in Sections 3.2 and 3.3 to determine (and star formation efficiencies (SFE), i.e. ) for the sources in our sample. Figure 12 shows our depletion time estimates versus redshift, along with the expected evolution for main-sequence star-forming galaxies and starburst type galaxies (Tacconi et al., 2018). Our are systematically shorter than those expected for main-sequence galaxies and consistent with the starburst population. These results, however, should be taken with caution since the are inferred from the ALMA 1.3 mm photometry and are not totally independent from the SED-based SFR determinations.
The inferred depletion times for the S - Fields cover a range from 0.2 Gyr to 0.4 Gyr. These values are also consistent with those found in the literature for similar samples, where the depletion times are between 0.1 Gyr and 1 Gyr (Hatsukade et al., 2015; Pantoni et al., 2021; Berta et al., 2023; Castillo et al., 2023), but lower compared to those found for nearby galaxies ( and Gyr, e.g Bigiel et al., 2008; Leroy et al., 2008; Genzel et al., 2010). If we scale our gas mass determinations to assume , our depletion times are reduced even further (30 to 60 Myr). These values are in agreement with those reported by the extreme SMG population reported to have high star formation efficiency (Aravena et al., 2016; Yang et al., 2017; Ikarashi et al., 2022).
Considering that per cent of the sources in our sample are individual (at 1 arcsecond resolution), we suggest that the processes that increase the SFE in these galaxies act at scales of kpc as are near-coalescence mergers and other physical processes such as gravitational instabilities and gas turbulence.
@@ 5 Summary and Conclusions
The median IR luminosities for the S - Fields and PLUM - Fields of and L⊙, correspond to SFR of and , respectively. This is in line with previous studies and confirms the extreme nature of these SMG, which also show high gas and dust content with and . Combining these measurements, we found average depletion times of Gyr, consistent with what has been measured for other SMG, implying high SFE compared to the main-sequence population. We stress, however, that these results should be taken with caution until a more direct measurement of the gas masses become available.
@@ Acknowledgements
We thank to the referee for insightful comments that have improved the manuscript. We also want to thank Olga Vega, Yalia Divakara Mayya, Manuel Zamora-Avilés, Elena Terlevich and Eric F. Jiménez-Andrade for their helpful discussions and comments, and to Michał Michałowski for insightful suggestions on the flux-deblending process. MQR would like to thank the Consejo Nacional de Humanidades Ciencias y Technologías (CONAHCYT) for her PhD grant. This work has been supported by the CONAHCYT through projects A1-S-45680 and CB 2016 - 281948. JAZ acknowledge funding from JSPS KAKENHI grant number KG23K13150. The Herschel-ATLAS is a project from Herschel, which is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. This paper use the following ALMA projects: ADS/JAO.ALMA#2016.1.00087.S, ADS/JAO.ALMA#2016.1.00139.S, ADS/JAO.ALMA#2017.1.00510.S, ADS/JAO.ALMA#2018.1.00489.S, ADS/JAO.ALMA#2018.1.00526.S, ADS/JAO.ALMA#2021.1.01628.S and ADSJAO.ALMA#2022.1.00432.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.